# Identifying the post-pandemic determinants of low performing students in Latin America through interpretable Machine Learning SHAP Values—Insights from PISA 2022

### v1

@@ 7 Conclusions
In this paper, I analyse the determinants of low performers students in the Latin American region at secondary level using PISA 2022 data, including 10 LAC countries. Combining the performance on three subjects assessed (math, reading and science), I focus on the bottom learning distribution into the three brackets of performance (OECD, 2023a): level 0 (no achievement), level 1 (below basic competencies) and level 2 (achieving basic competencies), where around 83% of the region students’ performance falls. I rely on two working samples: the sub-groups of bottom performers (level 0 versus level 1, $N$ = 16,236) and low performers (level 1 versus level 2, $N$ = 9,484) to compare the dealing determinants of only reaching either level 0 or level 1, correspondingly. Methodologically, using recent advancements on interpretable ML modelling (Chen et al., 2023; Lundberg and Lee, 2017), I use the SHAP method to pin down the main determinants of learning poverty in the LAC region. I obtain global and local versions of estimates, both at the regional level and country level.
The main findings from the global explanatory analysis can be summarised as follows. First, variables with the highest impacts (regardless of the direction of associations) are, to some extent, similar for determinants for bottom and low performers. Primary repetition, sense of belonging to school and well-being support from teachers are key drivers of falling into the lowest achievement group, whereas student’s paid work, the intensity of homework and mother’s educational level are relevant for the level 1 versus level 2 comparison (located higher in the learning distribution), with household wealth and educational support and digital devices being higher ranked than in the level 0-level 1 comparison. Second, the rate of school contextual poverty and STR (pedagogic support) are also more relevant for the reaching level 2, whilst the prevalence of students with a second language being more prominent as a driver for the level 0 performance group.
@@
Third, local estimations for students placed at the extremes of SHAP index distribution, indicate that a student with the highest probability of being a level 0 achiever attends a school with bad school climate, speak other minority language and had repeated, has no digital devices at home, comes from a poor family and works for payment half of the week. The school profile of this student shows wide disadvantages such as weak ICT infrastructure and teaching quality (only a third of teachers being certified), whereas a level 1 achiever attends a private with more autonomy and a low STR for pedagogic staff, more qualified teachers, and with larger connectivity (over 90% of PCs connected to internet). Similarly findings are obtained for the level 1 and level 2 comparison where —as well as a level 2 student being better off financially overall, and having more educational inputs and support—, he//she is not in the labour market having more feasible time for doing homework (twice as much as a level 1 student), and attends a school with considerable higher school’s rates for certification and PD attendance of teachers (80%-95%).
@@
Regarding countries’ estimates, I find quite homogenous patterns as far as global average contribution of top ranked factors is concerned, with repetition at primary, household wealth, and educational ICT inputs (digital devices) being top ten ranked covariates in at least 8 out of the 10 total countries. Other important covariates across countries are school safety, student work, books at home and mother’s education. There is, however, some between-country heterogeneity when it comes to local estimates and defining characteristics of a lowest achieving student. Level 0 (no achiever) shares, across countries, features such as being a repeater with very few devices, coming from a low SES family and attending a school with poor school climate, but with some differences and specific country drivers. In Chile, the lowest performer also skipped school and the education of his//her mother is below secondary while in Panama mothers’ education for this level 0 student profile is even worse (uncompleted primary); in Mexico, a new feature is that a level 0 student does paid-work for nearly the whole week (and missed school for 3 months); in Peru and Paraguay school features turn out to be prominent (very low ICT infrastructure, low autonomy and just 16% of PD teachers attendance in the case of Paraguay, and the lowest performing student for Peru is Indigenous). A Level 1 student profile is, in the majority countries, from a family with an above average level of wealth, has a fair number of digital devices, is not engaged in paid work, and mostly their schools they attend have a large degree of autonomy on their decisions.
@@ Appendix 0.B Shapley values and ML implementation
[…]
Shap values—estimation. When estimating a ML model, the aim is to explain a predictive model $f(x)$ trained on $\mathcal{X}=\{x^{[i]},y^{[i]}\}_{i=1}^{N_{\text{train}}}$, where $x^{[i]}$ is a feature vector of dimension $M$, $y^{[i]}$ is the vector of responses and $N_{\text{train}}$ is the number of observations used in the trained model. The prediction for a given feature vector value ($\widehat{y}=f(x=x^{*}$)) is obtained using Eq. (12) where the predictive model $f$ is the game, and the feature vector is the $M$ players. The $j$th feature in the prediction $f(x^{*})$ is given by the Shapley value $\phi_{j}$, where $f(x^{*})=\phi_{0}+\sum_{j=1}^{M}\phi^{*}_{j}$ and so the sum of Shapley values is the difference between the given prediction and the global average prediction (i.e., $\phi_{0}=\mathbb{E}[f(x)]$). To calculate Eq. (12), one needs to define a function which mirrors $f(x^{*})$ when only features on the coalition $\mathcal{S}$ are known: $\upsilon(\mathcal{S})=\upsilon(\mathcal{S},x^{*})$. This can be approximated by the contribution function (the expected response $f(x)$ conditioning on the features in $\mathcal{S}$ taking on the values $x_{\mathcal{S}}^{*}$):
$\upsilon(\mathcal{S})=\mathbb{E}\big[f(x_{\mathcal{\bar{S}}},x_{\mathcal{S}}|x_{\mathcal{S}}=x^{*}_{\mathcal{S}})\big]$ (13)
where $x_{\mathcal{\bar{S}}}$ denotes features outside the subset $\mathcal{S}$. The conditional expectation in Eq. (13) describes the whole probability distribution, and is a common estimator for predictions’ applications. Here, for the estimations of leading drivers of learning poverty, I follow Lundberg and Lee (2017) who introduced the SHAP (SHapley Additive exPlanation) method which unifies the class of additive feature attribution methods. There are various (interventional conditional expectation) algorithms to define the value function for a set of attributes (e.g., DeepSHAP, LinearSHAP, KernelSHAP, TreeSHAP).
In this paper, and due to computational issues and the fact that I use a stacked ML model (which is a combination of classification models), I rely on the LinearSHAP (Chen et al., 2020), the default Explainer of the SHAP Python package, which estimates the conditional expectation of $x$ given different subsets (for a model $f(x)=\beta X+b)$:
$\phi_{i}(f,x)$ $=$ $\frac{1}{M!}\sum_{R}\mathbb{E}\big[f(x)|x_{\mathcal{S}^{R}\cup i}\big]-\mathbb{E}\big[f(x)|x_{\mathcal{S}^{R}}\big]$ (14)
$=$ $\beta\frac{1}{M!}\sum_{R}\mathbb{E}\big[x|x_{\mathcal{S}^{R}\cup i}\big]-\mathbb{E}\big[x|x_{\mathcal{S}^{R}}\big]$ (15)
A key feature for the estimations which are used in the paper are SHAP feature importance plots, offering a summary of the contribution of each feature over a set of samples. The importance of each feature is calculated by taking the average of the absolute Shapley values per feature across all samples, using the following formula:
[…]
Table C1: Students and family variables description
@@
Table D1: Outcome (performance level 0 versus level 1). Comparison of ML models – parameters search
@@
  • (1) Within each learner//model, highest metrics are shown in bold. (2) Also, the chosen most efficient configuration of each model is highlighted in bold –that is, models with the largest number of top preforming metrics– and in case of a draw, I rely on the AUC metric (column 6) to reach a final selection.
Table D2: Outcome (performance level 0 versus level 1). Comparison of ML models – parameters search
@@
Logistic penalty = none 0.7305 0.8609 0.7591 0.4846 0.8068 0.7825
regression penalty = l2 0.7299 0.8548 0.7613 0.4943 0.8053 0.7823
Lasso C = 10 0.7533 0.8641 0.7816 0.5443 0.8208 0.8143
with CV penalty C = 50 0.7549 0.8702 0.7801 0.5373 0.8227 0.8139
C = 100 0.7527 0.8576 0.7843 0.5549 0.8193 0.8147
C = 200 0.7524 0.8571 0.7842 0.5549 0.8190 0.8147
Trees = 100, max depth = 10 0.7616 0.8571 0.7943 0.5812 0.8245 0.9305
Trees = 500, max depth = 5 0.7533 0.8483 0.7899 0.5742 0.8180 0.8083
Trees = 500, max depth = 10 0.7673 0.8609 0.7988 0.5909 0.8287 0.9319
Trees = 1000, max depth = 5 0.7530 0.8492 0.7890 0.5716 0.8180 0.8086
Trees = 1000, max depth = 10 0.7692 0.8627 0.7998 0.5926 0.8301 0.9330
  • (1) Within each learner//model, highest metrics are shown in bold. (2) Also, the chosen most efficient configuration of each model is highlighted in bold –that is, models with the largest number of top preforming metrics– and in case of a draw, I rely on the AUC metric (column 6) to reach a final selection.
Figure E1: Intersectionality—SHAP values interactions
@@
Figure E2: Model: level 1 versus level 2 (low performing group). Beeswarm dot plot and heatmap of individual SHAP values for each student//famly and school covariate across observations
@@
(a) Argentina
(b) Brazil
(c) Chile
(d) Colombia
(e) Dominican Rep.
(f) Mexico
(g) Panama
(h) Peru
(i) Paraguay
(j) Uruguay
Figure F1: Countries’ results for model: level 0 versus level 1 (bottom performing group). Student//famliy and school covariates average SHAP values (absolute values)

### v2

@@ 7 Conclusions
In this paper, we analyse the determinants of low performers students in the LAC region at secondary level using PISA 2022 data. Combining the performance on three subjects assessed (math, reading and science), we focus on the bottom learning distribution for three brackets of performance : level 0 (no achievement), level 1 (below basic competencies) and level 2 (achieving basic competencies), where around 83% of the region students’ performance falls. We rely on two working samples: the sub-groups of bottom performers (level 0 versus level 1, $N$ = 16,236) and low performers (level 1 versus level 2, $N$ = 9,484) to compare the determinants of only reaching either level 0 or level 1, correspondingly. Methodologically, we use the SHAP method (Chen et al., 2023; Lundberg and Lee, 2017) to pin down the main determinants of learning poverty in the LAC region. We obtain global and local versions of estimates, both at the regional and country levels.
@@
Leading findings can be summarised as follows. First–from the global analysis–variables with the highest impacts (regardless of the direction of associations) are, to a certain degree, similar for bottom and low performers: primary repetition, sense of belonging to school and well-being support from teachers are key drivers of falling into the lowest achievement group, whereas student’s paid work, the intensity of homework and mother’s educational level are relevant for the level 1 versus level 2 comparison (located higher in the learning distribution), with household wealth and educational support and digital devices being higher ranked than in the level 0-level 1 comparison. Second, the rate of school contextual poverty and STR (pedagogic support) are also more relevant for the reaching level 2 performance, whilst the prevalence of students with a second language being more prominent as a driver for the level 0 performance group. Third, local estimations looking for students placed at the extremes of SHAP index distribution, indicate that a student with the highest probability of being a level 0 achiever attends a school with bad school climate, speak a minority language and had repeated, has no digital devices at home, comes from a poor family and works for payment half of the week. The school profile of this student shows disadvantages such as weak ICT infrastructure and teaching quality (only a third of teachers being certified), whereas a level 1 achiever attends a private with more autonomy and a low STR for pedagogic staff, more qualified teachers and with larger connectivity (over 90% of PCs connected to internet). Similarly findings are obtained for the level 1 and level 2 comparison where –as well as a level 2 student being better off financially overall, and having more educational inputs and support–, he//she is not in the labour market having more feasible time for doing homework (twice as much as a level 1 student), and attends a school with considerable higher rates of teachers’ certification and PD attendance (80%-95%).
@@
Regarding countries’ estimates, we find quite homogeneous patterns as far as global average contribution of top ranked factors is concerned, with repetition at primary, household wealth, and educational ICT inputs (digital devices) being in top ten ranked covariates in at least 8 out of the 10 total countries. Other important covariates across countries are school safety, student work, books at home and mother’s education. There is, however, some between-country heterogeneity when it comes to local estimates and defining characteristics of a lowest achieving student. A level 0 (no achiever) student shares, across countries, features such as being a repeater with very few digital devices, comes from a low SES family and attends a school with poor school climate, but in some instances with some differences and specific countries’ drivers. In Chile, the lowest performer also skipped school and the education of his//her mother is below secondary while in Panama mothers’ education for this level 0 student profile is even worse (uncompleted primary); in Mexico, a new feature is that a level 0 student does paid-work for nearly the whole week (and missed school for 3 months); in Peru and Paraguay school features turn out to be prominent (very low ICT infrastructure, low autonomy and just 16% of PD teachers attendance in the case of Paraguay, and the lowest performing student for Peru is Indigenous). A level 1 student profile is, in the majority countries, from a family with an above average level of wealth, has a fair number of digital devices, is not engaged in paid work and, on the whole, attends schools with a large degree of autonomy on their functioning decisions.
@@ Appendix 0.A Evaluation metrics
Table C1: Outcome (performance level 0 versus level 1). Comparison of ML models – parameters search
@@
  • (1) Within each learner//model, highest metrics are shown in bold. (2) Also, the chosen most efficient configuration of each model is highlighted in bold (that is, models with the largest number of top preforming metrics) and in case of a draw, we rely on the AUC metric (column 6) to reach a final selection. (3) Some of the additional (default) parameters are as follows. Logitlasso: penalty(l1) solver(saga); GB: subsample(1), max_depth(3), max_features(None); RF: max_features(1.0), min_samples_split(2); NN: solver(adam), batch_size(auto).
Table C2: Outcome (performance level 1 versus level 2). Comparison of ML models – parameters search
@@
Logistic penalty = none 0.6383 0.7868 0.6630 0.4246 0.7196 0.7021
regression penalty = l2 0.6284 0.7947 0.6517 0.3891 0.7162 0.6941
Lasso C = 10 0.6486 0.7850 0.6735 0.4525 0.7250 0.7233
with CV penalty C = 50 0.6492 0.7877 0.6732 0.4499 0.7259 0.7233
C = 100 0.6476 0.8106 0.6652 0.4132 0.7307 0.7198
C = 200 0.6486 0.7938 0.6709 0.4398 0.7272 0.7231
Gradient boosting Trees = 100, learning rate = 0.01 0.6268 0.9507 0.6198 0.1610 0.7503 0.713
Trees = 100, learning rate = 0.10 0.6622 0.7912 0.6850 0.4766 0.7343 0.7868
Trees = 500, learning rate = 0.01 0.6512 0.8176 0.6667 0.4119 0.7345 0.7625
Trees = 500, learning rate = 0.10 0.6684 0.7692 0.6990 0.5234 0.7324 0.8647
Trees = 1000, learning rate = 0.01 0.6590 0.7956 0.6805 0.4626 0.7335 0.7863
Trees = 1000, learning rate = 0.10 0.6694 0.7595 0.7037 0.5399 0.7305 0.9057
Random forest Trees = 100, max depth = 5 0.6440 0.8564 0.6506 0.3384 0.7394 0.742
Trees = 100, max depth = 10 0.6481 0.7833 0.6735 0.4537 0.7242 0.9354
Trees = 500, max depth = 5 0.6429 0.8520 0.6507 0.3422 0.7379 0.7439
Trees = 500, max depth = 10 0.6533 0.7877 0.6773 0.4601 0.7283 0.9378
Trees = 1000, max depth = 5 0.6450 0.8520 0.6525 0.3473 0.7390 0.7432
Trees = 1000, max depth = 10 0.6585 0.7947 0.6802 0.4626 0.7330 0.9377
Neural net 1 hidden layer (200 nodes), activation = relu 0.6133 0.9568 0.6098 0.1191 0.7449 0.6999
1 hidden layer (200 nodes), activation = logistic 0.6429 0.6502 0.7179 0.6324 0.6824 0.7307
1 hidden layer (200 nodes), activation = tanh 0.6481 0.6485 0.7258 0.6477 0.6850 0.7502
3 hidden layers (200, 100 and 50 nodes), activation = relu 0.6263 0.8564 0.6361 0.2953 0.7300 0.7105
3 hidden layers (200, 100 and 50 nodes), activation = logistic 0.6502 0.7137 0.6995 0.5589 0.7065 0.7979
3 hidden layers (200, 100 and 50 nodes), activation = tanh 0.6388 0.7101 0.6877 0.5361 0.6987 0.7874
  • (1) Within each learner//model, highest metrics are shown in bold. (2) Also, the chosen most efficient configuration of each model is highlighted in bold (that is, models with the largest number of top preforming metrics) and in case of a draw, we rely on the AUC metric (column 6) to reach a final selection. (3) Some of the additional (default) parameters are as follows. Logitlasso: penalty(l1) solver(saga); GB: subsample(1), max_depth(3), max_features(None); RF: max_features(1.0), min_samples_split(2); NN: solver(adam), batch_size(auto).
Table C3: Comparison of different stacked models
ACC RC PR SP F1 AUC
(1) (2) (3) (4) (5) (6)
Panel A – Model Gradient boosting
Outcome – level 0 versus level 1
trees learning rates
100 0.01 0.7746 0.8688 0.8027 0.5970 0.8344 0.9191
100 0.10 0.7719 0.8795 0.7938 0.5689 0.8344 0.8792
500 0.01 0.7698 0.8725 0.7952 0.5759 0.8320 0.9125
500 0.10 0.7795 0.8776 0.8032 0.5944 0.8388 0.9145
1000 0.01 0.7746 0.8790 0.7970 0.5777 0.8360 0.8822
1000 0.10 0.7783 0.8734 0.8042 0.5988 0.8374 0.9299
Panel B – Lasso (with CV penalty)
Outcome – level 1 versus level 2
C
10 0.6741 0.7947 0.6960 0.5006 0.7421 0.8825
50 0.6741 0.7947 0.6960 0.5006 0.7421 0.8827
100 0.6752 0.8009 0.6950 0.4943 0.7442 0.8894
200 0.6736 0.7956 0.6952 0.4981 0.7420 0.8852
  • (1) For assessment of different stacked models we select the model with the highest weight from interaction 2 (see Table 4) for each outcome, and then calculate evaluation metrics for the stacked model using the parameters grid of each model. (2) Within each stacked model, highest metrics are shown in bold.
[…]
@@
Figure D2: Model: level 1 versus level 2 (low performing group). Beeswarm dot plot and heatmap of individual SHAP values for each student//famly and school covariate across observations
@@
Figure D3: Countries’ results for model: level 0 versus level 1 (bottom performing group). Student//famliy and school covariates average SHAP values (absolute values)
Appendix E. Analysis for the sub-sample of close down schools during the COVID-19 pandemic
Here, we look into the pandemic main determinants on students’ learning poverty chances for the sub-sample of schools which were shutdown (around 75%). We rely on the school questionnaire (from question 58) which gathers information only for the sub-sample of schools closed. For the level 0-level 1 comparison sample, the number of observations are reduced from 16,236 to 13,489 (with 668 out of the total 2,747 school remaining opened); and, for the level1-level2 comparison sample, this reduction is from 9,484 to 6,995 observations.
We assess the associations of the following covariates: school days of closure (SDC), school prevalence of students engaged in remote learning, and three different barriers for online learning (as well as their interactions with school days of closure) through dependence plots yielding ‘predicted probability ratios’. Specifically, partial dependence plots show the marginal effect of a given covariate on the prediction of the ML model. We obtain the predicted relative probability (RP) of the educational outcome measuring learning poverty (level 0 =1, level 1 = 1; and level1 = 1 and level 2 = 0). The RP of a given COVID-19 pandemic related covariate is defined as:
$\text{RP}=\frac{f_{\mathcal{S}}(x_{\mathcal{S}})}{(1/n)\sum_{i=1}^{n}f(x^{(i)})}$ (E.1)
which is the mean value of the stacked model predicted probability when fixing the specific covariate divided by the mean value of the model’s predicted probability, and $f_{\mathcal{S}}(x_{\mathcal{S}})$ is the partial function of the specific covariate. Results are displayed in Figure E1 (for bottom performers) and in Figure E2 (for low performers).
First, for the bottom performers comparison, on the one hand, Figure 1(a) shows how longer SDC (above the average of $\approx$ 270) pushes students chances towards the lowest bracket of performance (level 0). In particular, $\widehat{\text{RP}}$ are above one for the 400-700 SDC interval, reaching its peak at 500 days with a $\widehat{\text{RP}}$ = 1.08 where students attending schools closed by 500 days have 8% higher probability of falling into the level 0 performing group (instead of the level 1 achievement group). Besides, Figure 1(b) highlights the relevance of schools reaching significant and larger rates of students actually engaged in remote learning. Indeed, all estimated RP are above one if this rate is below 8 (where the percentage of students attended distance learning being between 61%-70%).
On the other hand, as far as barriers for students online learning during the pandemic is concerned (Figures 1(c) to 1(e)), RP estimates show an overall tendency where chances of students only reaching level 0 performance increases the higher the extent of barriers are. But, with $\widehat{\text{RP}}$ between 1.00-1.02 at given interval of barriers covariates, these effects are not as significant as the effects of SDC and a school’s rate of students involved remote learning. Furthermore, in the last three plots (i.e., Figures 1(f) to 1(h)) we interact each of these three barriers with SDC so that to capture whether the impact of barriers on level 0 performing likelihood is potentiated by longer periods of school closure. Estimated RP for the three variables reach maximum values of 1.02, 1.03 and 1.06, without an increasing tendency along higher values of barriers and higher values of SDC. Overall, the bottom performing analysis shown in the plots of Figure E1 conveys more relevance to the degree of school shutdowns and rates of remote learning rather than barriers for engagement and pedagogic continuity during the pandemic.
Figure E1: Model: level 0 versus level 1 (bottom performing group). Partial dependence plots showing predicted relative ratios (RP) for school days of closure, rate of engagement on remote learning and barriers for remote learning
Second, with regards to the low performing group comparison, pandemic covariates’ impact on the probability of students falling into level 1 (instead of level 2) has a distinct pattern when compared to what we found earlier for students chances falling into level 0 (shown in Figure E1). To begin with, SDC impacts on level 1 achievement is much smaller in magnitude, RP below 1.005 (Figures 2(a)), and for the variable school’s rate of engagement on remote learning of students (Figure 2(b)), there is a similar decreasing tendency showing the importance of students’ being engaged in distance learning for balancing level 1 and level 2 performance chances, though this become equal at a school’s higher rate for remote learning (when the covariate takes the value of 10: 81%-90%). In addition, barriers’ estimates (Figures 2(c)-Figure 2(e)) for distance learning have a clear increasing tendency (specially connectivity and systemic barriers) and, when they are above their mean values, the $\widehat{\text{RP}}$ are above 1, but again these values are not so large ($\leq$ 1.01). The same holds for the interacted versions of barriers (i.e., Figures 2(f) to Figure 2(h)).
Figure E2: Model: level 1 versus level 2 (low performing group). Partial dependence plots showing predicted relative ratios (RP) for school days of closure, rate of engagement on remote learning and barriers for remote learning